# Evolution of the Inner Accretion Flow in Swift J1727.8–1613 across Intermediate States: Insights from Broadband Spectral and Timing Analysis

### v1

@@ 4 Discussions
We have performed a broadband spectral study of the newly detected Galactic BHXRB Swift J1727.8–1613 during its 2023 outburst over the 0.72000.7-200 keV band in the HIMS using simultaneous NICER, NuSTAR, Insight-HXMT/HE, and AstroSat/CZTI observations, and over the 0.7780.7-78 keV band in the SIMS using contemporaneous observations from NICER and NuSTAR data. We also investigate the source variability across the two different spectral states. This outburst is found to be a complete one, similar to the canonical outbursts seen in low-mass BHXRBs. Substantial evolution in the spectral hardness also indicates a spectral state transition from the HIMS to the SIMS between the two observations.
4.1 Thermal Comptonization and Accretion Disk
@@
The analysis of the NuSTAR/FPMA spectral data reveals that the reflection component was notably weak during both observations (see Figure 3). Although residuals from both observations exhibit a distinct feature in the 676–7 keV band, indicative of an iron emission line, the characteristic reflection hump, typically peaking in the 204020–40 keV band, was either faint or absent. Additionally, an upward trend in the residuals above $\sim 50$ keV was observed during the HIMS but was not present during the SIMS. A similar high-energy excess above $40$ keV, along with a weak reflection hump, was previously reported in NuSTAR and Insight-HXMT observations during the LHS by Liu2024arXiv240603834L. The authors found the reflection hump in the Insight-HXMT data to be relatively prominent, peaking around $20$ keV. They proposed that this feature may be partially artificial, resulting from an improper modeling of the high-energy rollover by the cutoffpl model used in their spectral analysis.
@@
Our spectral analysis in the 0.7780.7–78 keV band during the HIMS reveals that the broadband spectrum requires models comprising thermal Comptonization of the disk photons in two physically distinct Comptonizing regions, their associated reflections, and thermal emission from the disk. The reflection due to the hard component originates from a distant part of the disk, where relativistic effects are weak, and strong illumination is observed. On the other hand, the reflection from the soft component, dominating at low energies, appears to be more blurred due to its proximity to the inner disk. Polarimetric studies of the source also favor radially extended coronal geometry confined to the disk plane, rather than being aligned along the jet axis (Ingram2024ApJ...968...76I; Veledina2023ApJ...958L..16V).
@@
We do not find any evidence for a high-energy tail, typically associated with Comptonization by non-thermal scattering electrons in the corona, up to 78 keV. It is worth mentioning here that Liu2024arXiv240603834L attributed the excess above $40$ keV observed in the LHS to the presence of a high-energy tail extending beyond the thermal Comptonization component. Moreover, Peng2024ApJ...960L..17P also suggested the excess above $40$ keV in the NuSTAR and Insight-HXMT spectral data taken in the transitioning period of the source towards the HIMS arises due to the presence of a hard X-ray tail coming from either a relativistic jet or from the base/corona below a slowly moving jet. The authors used a single Comptonizing region and a cutoff power-law model to account for the hard X-ray tail. However, a similar upward trend above $\sim 50$ keV (see Figure 3), observed in this work, is well reproduced by adopting a model that includes two thermal Comptonizing regions.
@@
Results from the reflection modeling suggest that the inner edge of the accretion disk is truncated well away from the ISCO. Although we are able to place a lower bound on the inner disk radius ($R_{\rm in}$), this parameter remains relatively insensitive to the specific choice of the continuum model, preventing a definitive constraint on the degree of truncation. The lack of a prominent reflection hump makes it difficult to robustly estimate $R_{\rm in}$ through reflection spectroscopy. However, an independent measurement of $R_{\rm in}\sim 29$$35~r_{\rm{g}}$ obtained from modeling the disk continuum also supports the scenario of a truncated accretion disk in the HIMS.
@@
The moderate to high disk ionization inferred from our analysis is likely due to strong coronal illumination. However, the relatively high intrinsic disk temperature ($\sim 0.4$ keV) may also contribute significantly to the elevated ionization state. The covering fraction ($C_{\rm{frac}}\sim 0.4$) suggests that a substantial portion of the disk photons are intercepted by the soft Comptonizing region, further supporting its interpretation as a warm, optically thick layer enveloping the disk. The overall spectral characteristics observed in the HIMS of Swift J1727.8–1613 closely resemble the broadband (0.71000.7–100 keV) spectrum of the low-mass BHXRB GX 339–4 in the same state (Chand2024ApJ...972...20C).
4.2 Hybrid Comptonization
@@
We notice that Comptonization by a purely thermal electron population fails to describe the high-energy excess at $\gtrsim 100$ keV (see Figure 5); instead, a hybrid electron distribution is required. In particular, $\gamma_{\rm{min}}\sim 1.3$, with electron power-law index of $2$, suggests that a significant portion of the high-energy emission arises from inverse Compton scattering of soft photons by the non-thermally distributed electrons. This provides strong evidence for a hybrid coronal plasma, where thermal electrons dominate the Comptonized spectrum at low energies, with a non-thermal tail emerging above a certain energy due to the non-thermally populated electrons. This hybrid model also favors the scenario where the disk is truncated far away from the ISCO, consistent with the presence of non-thermal particle acceleration within a hot and optically thin inner accretion flow.
@@
Evidence of the possible presence of a hard X-ray tail originating from non-thermal electrons in the LHS of Swift J1727.8–1613 during the same outburst was previously reported by Liu2024arXiv240603834L. Their results show that the contribution from the non-thermal electron population was relatively soft with a steeper distribution at the initial phase of the outburst, and the particle acceleration gradually increased as the source was transitioning towards the HIMS during the rising phase. This is consistent with our present results, where significant non-thermal particle acceleration is seen in the HIMS. Furthermore, hybrid Comptonization with a significant non-thermal electron population with $\gamma_{\rm{min}}~\sim 1.3$ and slightly steep electron power-law index of $\sim 2.5$ was also reported in the LHS of the bright low-mass BHXRB MAXI J1820+070 by Zdziarski2021ApJ...914L...5Z. This detection, along with a substantially truncated accretion disk in the LHS, was further confirmed in the follow-up study including Insight-HXMT data by Zdziarski2022ApJ...928...11Z.
@@
The electron temperature, $kT_{\rm{e}}\sim 13-18$ keV, of the hard Comptonizing region, inferred from all three spectral models (M1-M3) in the HIMS, remains low, which could be a consequence of the energy balance within the accretion flow. Alternatively, such a low $kT_{\rm e}$ could also result from efficient cooling by synchrotron photons produced by a hybrid electron distribution (Zdziarski2004PThPS.155...99Z; Malzac2009MNRAS.392..570M; Veledina2011ApJ...737L..17V; Zdziarski2021ApJ...914L...5Z).
4.3 State Transition and Disk Evolution
@@
In contrast to the HIMS, the coronal structure during Obs2 appears to be relatively simple and can be described with a single Comptonizing region. A relatively steeper photon index and a low covering fraction further indicate that the corona became weak in this state. We also notice a significant evolution in the accretion disk in terms of elevated inner disk temperature, ionization, and relatively closer extent (obtained from modeling both the reflection and disk continuum modeling) to the ISCO. These further suggest that the source transitioned to the SIMS during Obs2.
4.4 Spin, Mass, Distance, and Disk Inclination
@@
Reflection modeling alone (see model M4) could not constrain the spin properly, yielding a value $\gtrsim 0.7$ at the $90\%$ confidence limit. Peng2024ApJ...960L..17P reported a very high black hole spin of $0.98^{+0.02}_{-0.07}$ for Swift J1727.8–1613 by modeling the reflection continuum in the LHS. However, the reflection method for spin measurement relies on the assumption that the disk extends down to the ISCO or very close to it, and the extent of the inner accretion disk in the LHS is highly debated. On the other hand, our spin estimation of $0.79^{+0.03}_{-0.13}$ by modeling the disk continuum and reflection component together (see model M5), which provides a more physically self-consistent approach, is slightly lower but broadly consistent with the value reported by Svoboda2024ApJ...966L..35S, using the HSS observations of the source in February 2024.
@@
We find a close agreement between our estimated black hole mass, $M_{\rm{BH}}=10.5^{+7.7}_{-3.0}~M_{\odot}$ (see model M5), and previously reported values. This value also appears to be consistent with the black hole mass obtained from the mass function suggested by Mata2025A&A...693A.129M. Furthermore, zdziarski2025ApJ...986L..35Z favored a black hole mass of $\lesssim 13M_{\odot}$ for Swift J1727.8–1613, which aligns well with our estimate. Additionally, Debnath2025arXiv250416391D, using the propagating oscillatory shock model and the observed QPO–photon index correlation, reported a black hole mass of $\sim 13~M_{\odot}$ for this source that is also consistent with our results. Although the black hole mass inferred from bhspec appears to be consistent with that obtained from kerrbb within error bars, the slightly higher value is likely a consequence of the poorly constrained source distance in bhspec. It should be noted here that the bhspec model does not account for returning radiation, in which photons emitted from the inner disk are gravitationally bent back onto the disk surface. This process can introduce additional heating and alter the observed thermal spectrum, particularly in the high-energy tail. The effect becomes significant for systems with high black hole spin and high inclination. Although the spin and inclination inferred in this work represent a moderately relativistic regime, the omission of returning radiation may still introduce a minor systematic uncertainty in the derived disk parameters using this model. However, our estimate of the source distance, $D=3.5^{+1.3}_{-0.5}$ kpc, from the kerrbb component agrees well with the value reported by Mata2025A&A...693A.129M based on the optical spectroscopy.
@@
The disk inclination angle obtained from the models M5 and M6 lies in the range of $42^{\circ}$$50^{\circ}$, which is in good agreement with the $30^{\circ}$$50^{\circ}$ range estimated by Svoboda2024ApJ...966L..35S from polarimetric analysis of the HSS observation. Additionally, Peng2024ApJ...960L..17P also reported a disk inclination angle of $\sim 40^{\circ}$ using LHS observations in 2023. However, some studies reported the disk inclination angle in Swift J1727.8–1613 to be $\gtrsim 78^{\circ}$ (Chatterjee2024ApJ...977..148C; Debnath2024ApJ...975..194D), which is unlikely given the absence of any eclipse or dips in the source lightcurve.
4.5 Source Variability
@@
The shapes of the PDSs in different energy bands indicate a slight decrease in variability as the source transitioned to the SIMS. In both the spectral states, the increase in PDS power with energy suggests that high-energy coronal photons exhibit stronger variability than the softer ones, associated with the disk. In the HIMS, however, the PDSs above $30$ keV are nearly identical, suggesting a similar variability pattern among hard photons in this energy range. Additionally, the significant increase in QPO frequency from $\sim 1.3$ Hz to $\sim 6.6$ Hz suggests a change in the disk–corona geometry between the two states. If QPOs are thought to be associated with the inner edge of the disk or the outer boundary of the inner hot flow, the observed increase in the QPO frequency implies that the disk has moved relatively closer to the ISCO in the SIMS, potentially accompanied by a contraction of the corona. This interpretation is consistent with the results from our spectral modeling, where the disk is found to be moving closer to the ISCO in the SIMS.
@@
The shapes of the rms–energy spectra observed in the HIMS and SIMS differ from those reported by Gierlinski2005MNRAS.363.1349G for the LHS of XTE J1650–500 using RXTE data, where the variability is stronger at lower energies and gradually decreases with increasing photon energy. However, their study lacked coverage below $3$ keV, making it difficult to investigate the variability associated with the disk emission at softer energies. The morphology of the rms-energy spectra does not always remain the same and is found to vary across the spectral states and sources as well. In the intermediate states of the BHXRB XTE J1550–564, Gierlinski2005MNRAS.363.1349G reported that fractional rms increases with the increase in photon energy. This trend is consistent with our findings in both the HIMS and SIMS of Swift J1727.8–1613. In addition, a similar trend in the rms–energy spectrum derived at the QPO frequency was also reported during the HIMS of the low-mass BHXRB MAXI J1803–298 by Chand2022ApJ...933...69C, though their study also lacked the data below $3$ keV.
@@
Our investigation of the rms-energy spectra in both the spectral states reveals a weakly variable or stable disk and a highly variable power-law component, mainly governing the overall source variability. The nature of the variability remains similar between the two states, though a slight decrease is observed in the SIMS. Since our data in the HIMS is limited only up to $120$ keV, and our spectral modeling suggests that the contribution from non-thermal distribution becomes significant above $\sim 100$ keV, a robust statement on the variability associated with the non-thermal electron population is difficult. Moreover, the lack of high-energy data in the SIMS also restricts us from studying the variability pattern governed by the high-energy photons (>30>30 keV), and the non-thermal electron population in this state.
5 Conclusions

### v2

@@ IV Discussions
We have performed a broadband spectral study of the newly detected Galactic BHXRB Swift J1727.8–1613 during its 2023 outburst over the 0.72000.7-200 keV band in the HIMS using simultaneous NICER, NuSTAR, Insight-HXMT/HE, and AstroSat/CZTI observations, and over the 0.7780.7-78 keV band in the SIMS using contemporaneous observations from NICER and NuSTAR data. We also investigate source variability between the two spectral states. This outburst is complete, similar to the canonical outbursts observed in low-mass BHXRBs. Substantial evolution in spectral hardness also indicates a transition from the HIMS to the SIMS between the two observations.
IV.1 Thermal Comptonization and Accretion Disk
@@
The analysis of the NuSTAR/FPMA spectral data reveals that the reflection component was notably weak during both observations (see Figure 3). Although residuals from both observations exhibit a distinct feature in the 676–7 keV band, indicative of an iron emission line, the characteristic reflection hump, typically peaking in the 20–40 keV band, was either faint or absent. Additionally, an upward trend in the residuals above $\sim 50$ keV was observed during the HIMS but was not present during the SIMS. A similar high-energy excess above $40$ keV, along with a weak reflection hump, was previously reported in NuSTAR and Insight-HXMT observations during the LHS by Liu et al. (2024). The authors found the reflection hump in the Insight-HXMT data to be relatively prominent, peaking around 20 keV. They proposed that this feature may be partially artificial, resulting from an improper modeling of the high-energy rollover by the cutoffpl model used in their spectral analysis.
@@
Our spectral analysis in the 0.7780.7–78 keV band during HIMS indicates that the broadband spectrum requires models that include thermal Comptonization of disk photons in two physically distinct Comptonizing regions, their associated reflections, and thermal emission from the disk. The reflection from the hard component originates from a distant part of the disk, where relativistic effects are weak and strong illumination is observed. On the other hand, the reflection from the soft component, which dominates at low energies, appears more blurred due to its proximity to the inner disk. Polarimetric studies of the source also favor radially extended coronal geometry confined to the disk plane, rather than being aligned along the jet axis (Ingram et al., 2024; Veledina et al., 2023).
@@
We do not find any evidence for a high-energy tail, typically associated with Comptonization by non-thermal scattering electrons in the corona, up to 78 keV. It is worth mentioning here that Liu et al. (2024) attributed the excess above $40$ keV observed in the LHS to the presence of a high-energy tail extending beyond the thermal Comptonization component. Moreover, Peng et al. (2024) also suggested that the excess above $40$ keV in the NuSTAR and Insight-HXMT spectral data taken in the transitioning period of the source towards the HIMS arises due to the presence of a hard X-ray tail coming from either a relativistic jet or from the base/corona below a slowly moving jet. The authors used a single Comptonizing region and a cutoff power-law model to account for the hard X-ray tail. However, a similar upward trend above $\sim 50$ keV (see Figure 3) observed in this work is well reproduced by a model that includes two thermal Comptonizing regions.
@@
Results from the reflection modeling suggest that the inner edge of the accretion disk is truncated well away from the ISCO. Although we can place a lower bound on the inner disk radius ($R_{\rm in}$), this parameter remains relatively insensitive to the specific choice of the continuum model, preventing a definitive constraint on the degree of truncation. The absence of a prominent reflection hump makes it difficult to reliably estimate $R_{\rm in}$ using reflection spectroscopy. However, an independent measurement of $R_{\rm in}\sim 29$$35~r_{\rm{g}}$ obtained from modeling the disk continuum also supports the scenario of a truncated accretion disk in the HIMS.
@@
The moderate-to-high disk ionization inferred from our analysis is likely due to intense coronal illumination. However, the relatively high intrinsic disk temperature ($\sim 0.4$ keV) may also contribute significantly to the elevated ionization state. The covering fraction ($C_{\rm{frac}}\sim 0.4$) suggests that a substantial portion of the disk photons is intercepted by the soft Comptonizing region, further supporting its interpretation as a warm, optically thick layer enveloping the disk. The overall spectral characteristics observed in the HIMS of Swift J1727.8–1613 closely resemble the broadband (0.71000.7–100 keV) spectrum of the low-mass BHXRB GX 339–4 in the same state (Chand et al., 2024).
IV.2 Hybrid Comptonization
@@
We observe that Comptonization by a purely thermal electron population fails to account for the high-energy excess at $\gtrsim 100$ keV (see Figure 5); instead, a hybrid electron distribution is required. In particular, $\gamma_{\rm{min}}\sim 1.3$, with electron power-law index of $2$, suggests that a significant portion of the high-energy emission arises from inverse Compton scattering of soft photons by the non-thermally distributed electrons. This provides strong evidence for a hybrid coronal plasma, in which thermal electrons dominate the Comptonized spectrum at low energies, with a non-thermal tail emerging above a specific energy due to non-thermal electron populations. This hybrid model also favors the scenario in which the disk is truncated far from the ISCO, consistent with non-thermal particle acceleration in a hot, optically thin inner accretion flow.
@@
Evidence of the possible presence of a hard X-ray tail originating from non-thermal electrons in the LHS of Swift J1727.8–1613 during the same outburst was previously reported by Liu et al. (2024). Their results show that the contribution from the non-thermal electron population was relatively soft, with a steeper distribution at the initial phase of the outburst. That particle acceleration gradually increased as the source transitioned towards the HIMS during the rising phase. This is consistent with our current results, which show significant non-thermal particle acceleration in the HIMS. Furthermore, hybrid Comptonization with a significant non-thermal electron population with $\gamma_{\rm{min}}~\sim 1.3$ and slightly steep electron power-law index of $\sim 2.5$ was also reported in the LHS of the bright low-mass BHXRB MAXI J1820+070 by Zdziarski et al. (2021c). This detection, along with a substantially truncated accretion disk in the LHS, was further confirmed in a follow-up study that included Insight-HXMT data by Zdziarski et al. (2022).
@@
The electron temperature, $kT_{\rm{e}}\sim 13-18$ keV, of the hard Comptonizing region, inferred from all three spectral models (M1-M3) in the HIMS, remains low, which could be a consequence of the energy balance within the accretion flow. Alternatively, such a low $kT_{\rm e}$ could also result from efficient cooling by synchrotron photons produced by a hybrid electron distribution (Zdziarski and Gierliński, 2004; Malzac and Belmont, 2009; Veledina et al., 2011; Zdziarski et al., 2021c).
IV.3 State Transition and Disk Evolution
@@
In contrast to the HIMS, the coronal structure during Obs2 appears relatively simple and can be described by a single Comptonizing region. A relatively steeper photon index and a low covering fraction further indicate that the corona became weak in this state. We also observe a significant evolution in the accretion disk, with elevated inner-disk temperature and ionization, and a relatively closer extent to the ISCO (as inferred from reflection and disk-continuum modeling). These further suggest that the source transitioned to the SIMS during Obs2.
IV.4 Spin, Mass, Distance, and Disk Inclination
@@
Reflection modeling alone (see model M4) could not constrain the spin, yielding a value $\gtrsim 0.7$ at the $90\%$ confidence level. Peng et al. (2024) reported a very high black hole spin of $0.98^{+0.02}_{-0.07}$ for Swift J1727.8–1613 by modeling the reflection continuum in the LHS. However, the reflection method for spin measurement relies on the assumption that the disk extends down to the ISCO or very close to it, and the extent of the inner accretion disk in the LHS is highly debated. On the other hand, our spin estimation of $0.79^{+0.07}_{-0.15}$ by modeling the disk continuum and reflection component together (see model M5) is slightly lower but broadly consistent with the value reported by Svoboda et al. (2024), using the HSS observations of the source in February 2024.
@@
We find a close agreement between our estimated black hole mass, $M_{\rm{BH}}\approx 10~M_{\odot}$ (see model M5 and M6), and previously reported values. This value also appears to be consistent with the black hole mass obtained from the mass function suggested by Mata Sánchez et al. (2025). Furthermore, Zdziarski et al. (2025) favored a black hole mass of $\lesssim 13M_{\odot}$ for Swift J1727.8–1613, which aligns well with our estimate. Additionally, Debnath et al. (2025), using the propagating oscillatory shock model and the observed QPO–photon index correlation, reported a black hole mass of $\sim 13~M_{\odot}$ for this source that is also consistent with our results. It should be noted here that the bhspec model does not account for returning radiation, in which photons emitted from the inner disk are gravitationally bent back onto the disk surface. This process can introduce additional heating and alter the observed thermal spectrum, particularly in the high-energy tail. The effect becomes significant for systems with high black hole spin and high inclination. Thus, omission of the effect of returning radiation in the bhspec model may introduce a systematic uncertainty in the derived disk parameters.
@@
The disk inclination angle obtained from the models M5 and M6 lies in the range of $\sim 37\arcdeg$$53\arcdeg$, which is in good agreement with the $30\arcdeg$$50\arcdeg$ range estimated by Svoboda et al. (2024) from polarimetric analysis of the HSS observation. Additionally, Peng et al. (2024) also reported a disk inclination angle of $\sim 40\arcdeg$ using LHS observations in 2023. However, some studies reported the disk inclination angle in Swift J1727.8–1613 to be $\gtrsim 78\arcdeg$ (Chatterjee et al., 2024; Debnath et al., 2024), which is unlikely given the absence of any eclipse or dips in the source lightcurve.
IV.5 Source Variability
@@
The shapes of the PDSs in different energy bands indicate a slight decrease in variability as the source transitioned to the SIMS. In both spectral states, the increase in PDS power with energy suggests that high-energy coronal photons exhibit greater variability than the softer ones associated with the disk. In the HIMS, however, the PDSs above $30$ keV are nearly identical, suggesting a similar variability pattern among hard photons in this energy range. Additionally, the significant increase in QPO frequency from $\sim 1.3$ Hz to $\sim 6.6$ Hz suggests a change in the disk–corona geometry between the two states. If QPOs are thought to be associated with the inner edge of the disk or the outer boundary of the inner hot flow, the observed increase in the QPO frequency implies that the disk has moved relatively closer to the ISCO in the SIMS, potentially accompanied by a contraction of the corona. This interpretation is consistent with the results of our spectral modeling, which show that the disk moves closer to the ISCO in the SIMS.
@@
The shapes of the rms–energy spectra observed in the HIMS and SIMS differ from those reported by Gierliński and Zdziarski (2005) for the LHS of XTE J1650–500 using RXTE data, where the variability is more substantial at lower energies and gradually decreases with increasing photon energy. However, their study lacked coverage below $3$ keV, making it difficult to investigate the variability associated with the disk emission at softer energies. The morphology of the rms-energy spectra does not always remain constant and varies across spectral states and sources. In the intermediate states of the BHXRB XTE J1550–564, Gierliński and Zdziarski (2005) reported that fractional rms increases with the increase in photon energy. This trend is consistent with our findings from both HIMS and SIMS for Swift J1727.8–1613. In addition, a similar trend in the rms–energy spectrum derived at the QPO frequency was also reported during the HIMS of the low-mass BHXRB MAXI J1803–298 by Chand et al. (2022), though their study also lacked the data below $3$ keV.
@@
Our investigation of the rms-energy spectra in both spectral states reveals a weakly variable or stable disk and a highly variable power-law component, which primarily governs the overall source variability. The nature of the variability is similar across the two states, although a slight decrease is observed in the SIMS. Since our HIMS data are limited to $120$ keV, and our spectral modeling indicates that the contribution from the non-thermal distribution becomes significant above $\sim 100$ keV, a robust statement about the variability associated with the non-thermal electron population is difficult. Moreover, the lack of high-energy data in the SIMS also prevents us from studying the variability pattern driven by high-energy photons (>>30 keV) and the non-thermal electron population in this state.
V Conclusions